# Large Language Models Can Learn Temporal Reasoning

### v2

@@ 3.2.2 Data Augmentation
Figure 3: We have several strategies for data augmentation: remove irrelevant edges, use relation synonyms and change entities/times.
@@
Furthermore, to make sure LLM learns the underlying principle of symbolic temporal reasoning instead of just memorizing the training data, we introduce another two types of disturbances: (iii) globally map all the entity names in training data to some random names of the same type , and (iv) change the times based on a global offset, given in Eq. 7. For each relation, we generate these random names with GPT-3.5 by providing several examples of existing names.
$g^{\prime\prime}=f_{T}(f_{E}(g)),\ e^{\prime\prime}=f_{T}(e),$ (7)
$q^{\prime\prime}=f_{T}(f_{E}(q)),\ c^{\prime\prime}=f_{T}(f_{E}(c)),$
$a^{\prime\prime}=f_{T}(f_{E}(a))$
@@ 4 Experiments
We aim to answer the following research questions in our experiments: (1) Can we guide LLMs to perform more reliable symbolic reasoning with the strategies of bootstrapping CoTs and data augmentation? (2) Can our two-step framework lead to better temporal reasoning performance? (3) Does these learned capabilities of temporal reasoning generalize to other tasks?
@@ 4.1 Experimental Setup
In addition to TGQA, we consider the two existing datasets, TimeQA  (Chen et al., 2021) and TempReason  (Tan et al., 2023a), which are constructed using Wikipedia articles, excerpts, or summaries. Compared with these datasets, TGQA has a better quality due to the controlled generation, and more diverse question categories (see examples from other datasets in Appendix B). We evaluate our framework on all the datasets with the metrics of token-level F1, exact match (EM) and perplexity-based accuracy (Acc). Besides choosing F1 and EM, which are two basic metrics for span-based QA tasks, we consider Acc for LLM evaluations, i.e., selecting from a candidate set the answer with the lowest perplexity from LLMs as their prediction. (See rationales for evaluation metrics selection in Appendix B)
We compare our framework [^1] Code and data available at https://github.com/xiongsi heng/TG-LLM. with leading LLMs including Llama2 Touvron et al. (2023a), GPT-3.5 (Ouyang et al., 2022) and GPT-4 (OpenAI et al., 2023). We run the Llama2 family on local machines (4-bit quantization for Llama2-70B due to limited computational resources), and use APIs provided by OpenAI for GPT models. We evaluate their few-shot ICL performance on the test set. To avoid LLMs generating free text which affects the evaluation, we provide candidate answers for all questions. The model versions and prompt templates are provided in Appendix B and C, respectively. We also include the T5-based models  (Tan et al., 2023a; Yang et al., 2023a) for a comprehensive comparison. Results on TimeQA and TempReason reported in the original papers are used, and these models are fine-tuned and evaluated on TGQA.
@@ 4.2 Implementation Details
We use Llama2-13B as the baseline due to limited computational resources. We inject two adapters with selectors into the model for the text-to-TG translation and symbolic temporal reasoning. The adapters are trained in parallel. For inference, we first translate the original story into a temporal graph, and then perform symbolic reasoning on it, i.e., the adapters are used in sequence. In addition, we use GPT-3.5/GPT-4 for all the generation of story, graph and CoTs due to their high quality generation. All the prompt templates are given in Appendix C.
Figure 4: Performance comparison between different strategies of CoT generation on TGQA.
@@
Table 3: Human evaluation on the CoTs generated by different strategies. T1: using wrong info; T2: logical inconsistency; T3: external knowledge error; T4: temporal graph error. (See Appendix B for error type explanations)
@@
${}^{{\dagger}}$ 0.410 0.608 - 0.600 0.682 - 0.556 0.641 - 0.260 0.450 - 0.238 0.418 -
${}^{{\dagger}}$ 0.548 0.713 - 0.631 0.716 - 0.595 0.681 - 0.327 0.509 - 0.288 0.468 -
${}^{{\dagger}}$ 0.640 0.778 - - - - - - - 0.318 0.496 - 0.261 0.430 -
${}^{{\dagger}}$ 0.435 0.633 - 0.614 0.704 - 0.582 0.673 - 0.336 0.516 - 0.285 0.449 -
${}^{{\dagger}}$ 0.461 0.660 - 0.637 0.723 - 0.605 0.693 - 0.374 0.549 - 0.334 0.493 -
@@
Llama2-13B (SFT-TGSR) 0.797 0.850 0.819 0.664 0.691 0.673 0.631 0.664 0.649 0.424 0.522 0.432 0.356 0.469 0.399
Table 4: Main results using different models and strategies. We report exact match (EM), token-level F1 scores, and perplexity-based accuracy (Acc). Note: (1) Results with ${}^{{\dagger}}$ are reported in the original papers. We fine-tune and evaluate the models on our dataset. (2) Results with * are evaluated on 1000 random test samples.
Figure 5: Performance comparison between different strategies on TimeQA and TempReason. To obtain a fair comparison, we use Llama2-13B as the base model for all strategies. The basic strategy is in-context learning with standard Input/Output (ICL-SP), which is used to calculate the performance changes.
@@ 4.3 Main Results
Can our strategy bring more reliable symbolic reasoning?   Figure 4 shows the comparison between ICL with CoTs, SFT with CoTs, bootstrapping CoTs, and data augmentation on TGQA. It can be seen that LLMs learn better temporal reasoning with SFT than ICL. By providing CoTs with bootstrapping and our data augmentation strategies, the model performance gets further enhanced. Furthermore, inspired by  (Wang et al., 2023), we manually check 100 CoTs generated by different strategies. Evaluators are asked to classify the errors into four types (using wrong info, logical inconsistency, external knowledge error, and temporal graph error). It can be seen in Table 3 that our strategies reduce all types of error rates (ER).
Can our two-step framework lead to better temporal reasoning performance?   Table 4 shows the comparison between different models on all datasets. First, we observed that among all the LLMs with ICL, GPT-4 has the strongest performance as expected. For the Llama2 family, larger models have better performance due to advanced context understanding and improved generalization. We also found that CoTs not always bring better temporal reasoning due to unreliable intermediate results and hallucinations. More importantly, our model (SFT-TGSR), which is based on Llama2-13B, shows a comparable or even better performance than GPT-4 on all datasets. To prove the effectiveness of our framework, we also provide the results of some alternative strategies, where SFT-SP and SFT-CoT denote supervised fine-tuning on standard Input/Output prompts and CoTs, respectively. It can be seen that the two-step framework brings a substantial improvement in model performance on different datasets. We hypothesize that this improvement in performance is because our two-step neuro-symbolic process provides an easier path toward answering the temporal questions for LLMs.
Does these learned capabilities of temporal reasoning generalize to other tasks?  Figure 5 shows the comparison between different strategies on TimeQA and TempReason. In our framework, LLMs can learn the capability of text-to-graph translation and symbolic reasoning, which brings better temporal reasoning. More importantly, we observe that per-training on TGQA improves the model performance compared with ICL. Also, it can be concluded that these necessary capabilities in temporal reasoning are generalizable to different data distributions. Since the synthetic dataset is fully controllable and requires minimal supervision, we actually provide a way of improving the temporal reasoning capability of LLMs.
@@ 4.4 Ablation Study
We ablate different modules to see their contributions to the performance. Table 5 shows the performance comparison between different configurations. To obtain a fair comparison, we use Llama2-13B as the base model for all configurations. From the ablation study, we obtain some insights: (1) LLMs can benefit from explicitly presented (temporal) graph which is intuitive, concise and structured. (2) Given a reliable graph, symbolic reasoning brings better performance than free-text CoTs. (3) Data augmentation is necessary for LLMs to perform complex tasks such as temporal reasoning. (4) The introduction of external knowledge such as mathematics and commonsense can augment the generation.
@@ 5 Related Work
Language-based Temporal Reasoning  Recently, language-based TR has gained substantial research focus  (Nylund et al., 2023; Hu et al., 2023; Gurnee and Tegmark, 2023; Liu et al., 2023). The vision here is to help LMs understand temporal concepts. Existing methods mainly solve this problem with time-aware modeling. For example,  (Rosin et al., 2022; Pereira, 2022; Tan et al., 2023a) propose pre-training/fine-tuning strategies for robust TR.  (Ning et al., 2019; Zhou et al., 2020; Yang et al., 2020; Zhou et al., 2021; Yang et al., 2023a) design auxiliary objectives to introduce external knowledge. Although these methods make some progress, the underlying structure and logic of TR are either ignored or not explicitly involved.
Reasoning towards LMs   Although LLMs exhibit some emergent behaviors, it is still unknown whether they can actually perform reasoning and how strong their capability of reasoning is (Huang and Chang, 2023). Existing methods that try to elicit or improve reasoning can be divided into two directions: reasoning-involved modeling or hybrid methods. For example,  (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2023) use prompt engineering where some examples including intermediate reasoning steps are provided in context.  (Talmor et al., 2020; Nye et al., 2021; Wang et al., 2023; Ho et al., 2023) fine-tune LMs with these information before producing final answers. On the other hand,  (Li et al., 2023; Kynoch et al., 2023; Zhuang et al., 2023; Xu et al., 2023) combine LMs with domain-specific tools, empowering the model to perform more complex tasks that require reasoning.
Neuro-symbolic Reasoning Recent research has shown the advantages of combining neural networks with symbolic reasoning  (Sarker et al., 2021; Hamilton et al., 2022). In particular, knowledge graphs (KGs) as the foundation of (temporal) symbolic reasoning has been widely adopted in the past (Yang et al., 2017; Payani and Fekri, 2019a, b; Yang and Song, 2020; Qu et al., 2021). Moreover,  (Liu et al., 2022; Xiong et al., 2022, 2023) develop approaches for symbolic reasoning over temporal KGs. In this paper, we build the connection between language-based and temporal KG-based temporal symbolic reasoning. This connection brings the potential for extending these KG-based methods to language-based tasks. Different from existing methods (Luo et al., 2023; Zhang et al., 2023; Yuan et al., 2023), which limit their application to certain datasets, our framework offers enhanced generalization and usability, largely attributed to its innovative use of text-to-graph translation as a precursor to symbolic reasoning.
@@
TG + SR(bs) 0.723 0.797 0.736
TG + SR(bs + aug) 0.782 0.838 0.795
TG + EK + SR(bs) 0.753 0.803 0.776
TG + EK + SR(bs + aug) 0.797 0.850 0.819
Table 5: Ablation study results on TGQA. To obtain a fair comparison, we use Llama2-13B as the base model for all configurations. TG: temporal graph, SR: symbolic reasoning, EK: external knwoledge, bs: bootstrapping, aug: data augmentation.
@@ 6 Conclusion
TG-LLM, a new framework for language models, has been proposed to improve their performance on temporal reasoning. Besides producing the final answers, our framework also provides the temporal graph and symbolic reasoning steps. Experiments indicate that TG-LLM achieves better performance than existing methods. An interesting direction for future work is to extend it to more complex applications such as induction and abductive reasoning. Due to the graph structure and capability of symbolic reasoning, it is promising to improve the performance of LLMs on these tasks.
@@
Graph-augmented approaches such as TG-LLM help language models better learn related concepts from the perspective of graph. Although we demonstrate TG-LLM has good performance on understanding temporal relations, it still needs adaptations for temporal commonsense reasoning  (Zhou et al., 2019; Qin et al., 2021). Explicit in-context integration of commonsense presents opportunities for this task. Further, we mainly improve the capability of LLMs by introducing a new paradigm and providing more plausible and informative training data. There can be opportunities such as simulating an environment to provide feedback to LLMs  (Yao et al., 2023). For example, we can verify the generated TG based on prior knowledge such as the time gap between someone’s birth date and death date, i.e., a person’s lifespan, should fall into a certain range. In this way, we can further improve the quality of generation.
@@
Appendix A TGQA Dataset Statistics
In this section, we show statistics for our TGQA dataset. Figure 6 shows the number of facts distribution in training and test set, which determine the complexity of the task to some extent. Figure 7 shows the types of questions distribution in training and test set, where Q0: "Which event occurred first, <Event_A> or <Event_B>?", Q1: "Given the following <N> events: <Event_A>, <Event_B>, <Event_C>, <Event_D>, $\cdots$, which event is the first/second/third/fourth/$\cdots$ one in the chronological order?", Q2: "How long did the event <Event_A> last?", Q3: "", Q4: "How much time passed between the start of <Event_A> and the start of <Event_B>?", Q5: "What happened right before/after <Event_A> started?", Q6: "When did the <Event_A> occur?", Q7: "True or false: <Event_A> and <Event_B> happened at the same year?", Q8: "True or false: <Event_A> was still happening when <Event_B> started?". Note that we give eight categories of question in Table 2, since in a strict sense Q0 and Q1 can be considered as the same type. Among all the question types, Q1 has the largest portion since it has multiple variants. Similarly, Q7 with two variants also has a larger portion. To mitigate question category imbalance, we first calculate the metrics for each category, and then use the average as final scores. We also provide examples for the global mapping for entity names (Table 6) and external knowledge (Table 7) used in our dataset.

### v3

@@ 3.2.2 Graph Data Augmentation
Figure 3: We have several strategies for graph data augmentation: remove irrelevant edges, use relation synonyms and change entities/times.
@@
Furthermore, to make sure LLM learns the underlying logic of TR instead of just memorizing semantic information, we introduce another two types of disturbances: (iii) globally map all the entity names in training data to some random names of the same type (Eq.7) , and (iv) change the times based on a global offset (Eq.7). For each relation, we generate these random entity names with GPT-3.5 by providing several examples of existing names.
$g^{\prime}=f_{T}(f_{E}(g)),\ e^{\prime}=f_{T}(e),$ (7)
$q^{\prime}=f_{T}(f_{E}(q)),\ c^{\prime}=f_{T}(f_{E}(c)),$
$a^{\prime}=f_{T}(f_{E}(a))$
@@ 4 Experiments
We aim to answer the following research questions in our experiments: (1) Can our strategies (CoT bootstrapping and graph data augmentation) bring more reliable reasoning over TGs? (2) Can our two-step framework lead to better TR performance? (3) Do these learned capabilities of TR generalize to other tasks?
@@ 4.1 Experimental Setup
We demonstrate TG-LLM is a general framework by applying it to, besides TGQA, the two existing datasets, TimeQA  (Chen et al., 2021) and TempReason  (Tan et al., 2023a), which are constructed using Wikipedia articles, excerpts, or summaries. Examples from other datasets are listed in Appendix B. We thoroughly evaluate our framework on all the datasets with a combination of the metrics, i.e., token-level F1, exact match (EM) and perplexity-based accuracy (Acc). Besides choosing F1 and EM, which are two basic metrics for span-based QA tasks, we consider Acc for LLM evaluation, i.e., selecting from a candidate set the final answer with the lowest perplexity as the prediction. The construction of the candidate set for all datasets and detailed explanations are listed in Appendix B.
We primarily compare our framework with the leading LLMs, i.e., Llama2 Touvron et al. (2023a), GPT-3.5 (Ouyang et al., 2022) and GPT-4 (OpenAI et al., 2023). Specifically, we ran the Llama2 family on local machines (4-bit quantization for Llama2-70B due to limited computational resources), and used APIs provided by OpenAI for GPT models. We evaluated their few-shot ICL performance on the test set. To prove the effectiveness of our method, we also show the model performance with SFT on standard Input/Output prompts (SP) and CoTs from GPT-3.5. The model versions and prompt templates are provided in Appendix B and C, respectively. We also include the T5-based models  (Tan et al., 2023a; Yang et al., 2023a) for a comprehensive comparison. Results on TimeQA and TempReason reported in the original papers are used, and these models are fine-tuned and evaluated on TGQA.
@@ 4.2 Implementation Details
We use Llama2-13B as the baseline due to limited computational resources. We inject two adapters with selectors into the base model for the text-to-TG translation and temporal knowledge graph reasoning. The adapters are trained in parallel. For inference, we first translate the original story into a temporal graph, and then perform reasoning on it, i.e., the adapters are used in sequence. For data generation, we use GPT-3.5 for story, TG and CoT generation, and the verification of stories and TGs. We use GPT-4 to create the ICL demonstrations of CoT generation, due to its high generation quality. All the prompt templates are given in Appendix C.
Figure 4: Performance comparison between different CoT generation strategies on TGQA.
@@
Table 3: Human evaluation on the generated CoTs by different strategies. ER: error rate; T1: using wrong info; T2: logical inconsistency; T3: external knowledge error; T4: temporal graph error. (Error type explanations are listed in Appendix B.)
@@
T5-base 0.410 0.608 - 0.600 0.682 - 0.556 0.641 - 0.260 0.450 - 0.238 0.418 -
T5-large 0.548 0.713 - 0.631 0.716 - 0.595 0.681 - 0.327 0.509 - 0.288 0.468 -
Temp-T5 0.640 0.778 - - - - - - - 0.318 0.496 - 0.261 0.430 -
REMEMO-base 0.435 0.633 - 0.614 0.704 - 0.582 0.673 - 0.336 0.516 - 0.285 0.449 -
REMEMO-large 0.461 0.660 - 0.637 0.723 - 0.605 0.693 - 0.374 0.549 - 0.334 0.493 -
@@
Llama2-13B (SFT-TGR) 0.797 0.850 0.819 0.664 0.691 0.673 0.631 0.664 0.649 0.424 0.522 0.432 0.356 0.469 0.399
Table 4: Main results using different models and strategies. We report exact match (EM), token-level F1 scores, and perplexity-based accuracy (Acc). Note: (1) Results with are reported in the original papers. We only fine-tune and evaluate the models on our dataset. (2) Results with * are evaluated on 1000 random test samples.
Figure 5: Performance comparison between different strategies on TimeQA and TempReason. To obtain a fair comparison, we use Llama2-13B as the base model for all strategies. The basic strategy used to calculate the performance changes is in-context learning with standard Input/Output prompt (ICL-SP).
@@ 4.3 Main Results
Can our strategies bring more reliable reasoning over TGs? We show the comparison between ICL with CoTs, SFT with CoTs, bootstrapping CoTs, and graph data augmentation on TGQA (Figure 4). It can be seen that LLM learns TR better with SFT than ICL. By providing CoTs with bootstrapping and graph data augmentation strategies, the model performance gets further enhanced. Furthermore, inspired by  (Wang et al., 2023), we manually check 100 CoTs generated by different strategies (Table 3). Evaluators are asked to classify the errors into four types (using wrong info, logical inconsistency, external knowledge error, and temporal graph error). It can be seen that our strategies reduce all types of error rate.
Can our two-step framework lead to better temporal reasoning performance? We show the comparison between different models and strategies on all datasets (Table 4). First, we observed that among all the LLMs with ICL, GPT-4 has the strongest performance as expected. For the Llama2 family, larger models have better performance due to advanced context understanding and improved generalization. We also found that CoTs not always bring better TR due to unreliable intermediate results and hallucinations. From the results of some alternative strategies, where SFT-SP and SFT-CoT denote supervised fine-tuning on standard Input/Output prompts and vanilla CoT distillation, respectively, we prove the effectiveness of our framework (SFT-TGR). More importantly, our model, which is based on Llama2-13B, shows a comparable or even better performance than GPT-4 on all datasets. It can be seen that the two-step framework brings a substantial performance improvement on different datasets. We hypothesize this performance improvement is because our two-step reasoning process provides an easier path toward answering the temporal questions for LLM.
Do these learned capabilities of temporal reasoning generalize to other tasks? We show the comparison between different strategies (ICL with SP/CoT, SFT with TGQA/original data) on the two existing datasets, TimeQA and TempReason (Figure 5). Our framework learns the capabilities of text-to-TG translation and temporal graph reasoning, which brings better TR. More importantly, we observed that SFT on TGQA improves the model performance compared with ICL. It can be concluded that these necessary capabilities in TR are generalizable to different data distributions. Since TGQA is fully controllable and requires minimal supervision, we actually provide a general and effective way of TR capability improvement.
@@ 4.4 Ablation Study
We ablate different modules to see their contributions to the performance. We show the performance comparison between different configurations (Table 5). To obtain a fair comparison, we use Llama2-13B as the base model for all configurations. From the ablation study, we obtain some insights: (1) LLM can benefit from explicitly presented (temporal) graph which is intuitive, concise and structured. (2) Given a reliable graph, CoT bootstrapping with contrastive learning brings better performance than vanilla CoT distillation. (3) Data augmentation is necessary for LLMs to perform complex tasks such as temporal reasoning. (4) The introduction of external knowledge such as mathematics and commonsense can further augment the generation.
@@ 5 Related Work
Language-based Temporal Reasoning. Recently, language-based TR has gained substantial research focus  (Nylund et al., 2023; Hu et al., 2023; Gurnee and Tegmark, 2023; Liu et al., 2023; Vashishtha et al., 2020). The vision here is to help LMs understand temporal concepts. Existing methods mainly solve this problem with time-aware modeling. For example,  (Rosin et al., 2022; Pereira, 2022; Tan et al., 2023a) propose pre-training/fine-tuning strategies for robust TR.  (Ning et al., 2019; Zhou et al., 2020; Yang et al., 2020; Zhou et al., 2021; Yang et al., 2023a) design auxiliary objectives to introduce external knowledge. Although these methods make some progress, the underlying structure and logic of TR are either ignored or not explicitly involved.
Reasoning towards LMs. Although LLMs exhibit some emergent behaviors (Zhang et al., 2023a, 2024; Lai et al., 2024a, b; Wang et al., 2024; Chen et al., 2024a), it is still unknown whether they can actually perform reasoning and how strong their capability of reasoning is (Huang and Chang, 2023). Existing methods that try to elicit or improve reasoning can be divided into two directions: reasoning-involved modeling or hybrid methods. For example,  (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2023) use prompt engineering where some examples including intermediate reasoning steps are provided in context.  (Talmor et al., 2020; Nye et al., 2021; Wang et al., 2023; Ho et al., 2023) fine-tune LMs with these information before producing final answers. On the other hand,  (Li et al., 2023; Kynoch et al., 2023; Zhuang et al., 2023; Xu et al., 2023) combine LMs with domain-specific tools, empowering the model to perform more complex tasks that require reasoning.
Reasoning over Knowledge Graphs. Knowledge graphs (KGs) as the foundation representation of semantic and symbolic reasoning have been widely adopted in the past (Yang et al., 2017; Sun et al., 2018; Payani and Fekri, 2019a, b; Yang and Song, 2020; Qu et al., 2021; Huang et al., 2023; Chen et al., 2023, 2024b; Wan et al., 2024; Li et al., 2024). Moreover,  (Liu et al., 2022; Yang et al., 2022; Xiong et al., 2022, 2023) develop approaches for symbolic reasoning over temporal KGs. In this paper, we build the connection between language-based and symbolic-based temporal reasoning. This connection brings the potential for extending these KG-based methods to language-based tasks. Different from existing methods (Luo et al., 2023; Zhang et al., 2023b; Yuan et al., 2023), which limit their application to other tasks, our framework offers enhanced generalization and usability, largely attributed to its innovative use of text-to-graph translation as a precursor to deliberate reasoning.
@@
TG + CoT(bs) 0.723 0.797 0.736
TG + CoT(bs + aug) 0.782 0.838 0.795
TG + EK + CoT(bs) 0.753 0.803 0.776
TG + EK + CoT(bs + aug) 0.797 0.850 0.819
Table 5: Ablation study results on TGQA with Llama2-13B. TG: temporal graph, CoT: chain of thought, EK: external knwoledge, bs: bootstrapping, aug: graph data augmentation.
@@ 6 Conclusion
TG-LLM, a novel framework for LMs, has been proposed to improve their performance on temporal reasoning. To produce reliable final answers, our framework equips LLMs with the temporal graph and intermediate reasoning steps. Extensive experiments indicate that TG-LLM achieves better performance than existing pipelines. An interesting direction for future work is to extend it to more complex applications such as inductive and abductive reasoning. Due to the graph structure and capability of deliberate reasoning, it is promising to improve the model performance on these tasks.
@@
Graph-augmented approaches such as TG-LLM help language models better learn related concepts from the perspective of graph. Although we demonstrate TG-LLM has good performance on understanding temporal relations, it still needs adaptations for temporal commonsense reasoning  (Zhou et al., 2019; Qin et al., 2021). Explicit in-context integration of commonsense presents opportunities for this task. Further, we mainly improve the capability of LLMs by introducing a new paradigm and providing more plausible and informative training data. There can be opportunities such as simulating an environment to provide feedback to LLMs  (Yao et al., 2023). For example, we can verify the generated TG based on prior knowledge such as the time gap between someone’s birth date and death date, i.e., a person’s lifespan, should fall into a certain range. In this way, we can further improve the quality of generated intermediate steps.
@@
Appendix A Dataset Statistics of TGQA
In this section, we provide statistics for our TGQA dataset. We obtain 400 samples for training, 100 for validation and another 100 for test, with about 30 QA pairs in a single sample. We show the distribution of the number of facts (in one sample) in training (with validation) and test set (Figure 6). It determines the complexity of the TR tasks to some extent. We also show the distribution of question types in training (with validation) and test set (Figure 7), where Q0: "Which event occurred first, <Event_A> or <Event_B>?", Q1: "Given the following <N> events: <Event_A>, <Event_B>, <Event_C>, <Event_D>, $\cdots$, which event is the first/second/third/fourth/$\cdots$ one in the chronological order?", Q2: "How long did the event <Event_A> last?", Q3: "", Q4: "How much time passed between the start of <Event_A> and the start of <Event_B>?", Q5: "What happened right before/after <Event_A> started?", Q6: "When did the <Event_A> occur?", Q7: "True or false: <Event_A> and <Event_B> happened at the same year?", Q8: "True or false: <Event_A> was still happening when <Event_B> started?". Note that we give eight question categories in Table 2, since in a strict sense Q0 and Q1 can be considered as the same type. Among all the question types, Q1 has the largest portion since it has multiple variants. Similarly, Q7 with two variants also has a larger portion. To mitigate question category imbalance, we first calculate the metrics for each category, and then use the average as the final scores. Additionally, we show examples for the global mapping for entity names (Table 6) and external knowledge (Table 7) used in our dataset.
Figure 6: Number of facts distribution in TGQA.
Figure 7: Types of questions distribution in TGQA.
Appendix B Experiment Details
In this section, we present more experiment details for further research. We include graph data augmentation examples, model versions and evaluation tasks & metrics.